# Horizon Entropy Refined: Quantum Contributions and Cosmological Insights

### v1

@@ II Cosmology of Holographic Entropies
However, the crucial issue is related to resolving the crisis faced by the standard model of cosmology. In this direction, akin to modified gravity theories, modifying the entropy associated with the horizon also influences the equations governing the evolution of the universe Nojiri et al. (2021). In fact, there are many proposals for generalized versions of the entropy. Taking into account systems with long-range interactions, such as gravitational systems, which are categorized as non-extensive systems, Tsallis introduced a generalization of standard thermodynamics in 1988 Tsallis (1988). This extension broadens its applicability to non-extensive scenarios while retaining the standard Boltzmann-Gibbs theory as a limiting case Tsallis (1988). It is, thus, noteworthy that, when equipped with a specific entropy expression within any gravity theory, it becomes possible to rephrase the Friedmann equations of the Friedmann-Robertson-Walker (FRW) universe, which are extracted from the first law of thermodynamics on the apparent horizon, and vice versa. Considering this procedure, recent developments have led to significant interest in utilizing non-extensive Tsallis entropy for the entropy of gravitational systems Saridakis et al. (2018); Tavayef et al. (2018); da Silva and Silva (2021). This interest extends to the thermodynamic interpretation of gravitational field equations and the consequential effects of the non-extensive parameter within cosmology. Notably, this can result in variations in the gravitational constant’s strength, which consequently impacts the energy density of the dark components of the universe. This modification can explain the late-time acceleration of the universe without invoking a dark energy component.
@@ III Generalized Entropy in Our Model
Here we consider a different approach to suggest an alternative form of entropy, which can be motivated from the quantum gravity perspective.
Figure 1: Surface of a black hole with exaggerated roughness at the Planck scale resolution, illustrating quantum fluctuations and the fuzziness of the event horizon
According to general relativity, the horizon of a black hole is precisely defined by the Schwarzschild radius. On the other hand, from a quantum mechanical perspective, understanding the geometry of spacetime requires measurements using tools such as rods and clocks at different points, each subject to inherent uncertainties. This principle implies that any prospective quantum theory of gravity, built upon the foundations of quantum mechanics, must respect these uncertainty principles.
Therefore, considering the limitations on measurement precision imposed by quantum mechanics, it is reasonable to assume a degree of roughness in the area of a black hole when examining its surface at the Planck scale, as illustrated in Fig. 1. This roughness manifests as a random distribution, potentially rooted in a quantum theory of gravity where uncertainty plays a pivotal role. Evidently, larger surface areas correspond to a greater number of fluctuations, while smaller surfaces exhibit fewer fluctuations.
Assuming that these fluctuations are homogeneously distributed, the number of fluctuations is expected to scale proportionally with the surface area. For instance, doubling the surface area would result in a doubling of the number of fluctuations. Based on this linear relationship, a simple estimate for the corrected area can be expressed as
@@
where $A_{0}$ represents the area of the black hole corresponding to its Schwarzschild radius, and the second term accounts for the roughness term. Here, $\gamma$ serves as a parameter quantifying the quantum mechanical fuzziness in the black hole’s surface area. To approach the problem from a different perspective that may help enhance our physical intuition, we assume that the black hole surface consists of small pixels of size $l$. We choose this size to be very small so that summing these pixels yields a more accurate representation of the black hole surface. In other words, each pixel contains a few of the smallest possible pixels of size $l_{p}$. Fluctuations in the radius introduce can be of both positive and negative curvature distortions on the surface, each with the size $l_{p}$. Consequently, we can assume that the area of a pixel affected by quantum uncertainty in measurements can be expressed as
$d\mathcal{A}=l^{2}+\eta l_{p}^{2},$ (3)
where $\eta$ is a constant related to the probability distribution of quantum fluctuations. Regardless of whether the surface experiences positive or negative distortion, the area of the selected portion increases by a factor proportional to the Planck area, defined as $A_{p}=l_{p}^{2}$. The larger pixel size is also on the order of a few times the Planck scale (i.e., $l\propto l_{p}$), we can express the correction terms for the black hole’s area as in Eq. 2. It is worth mentioning that, although the volume on average doesn’t change, tiny ripples have the potential to increase the surface by several orders of magnitude. Considering these concepts, we can conceptualize the general form of the black hole entropy as
@@
In this model, $A_{0}$ once again denotes the horizon area. To test the validity of this model, we explore its cosmological consequences using the thermodynamic interpretation of the gravitational field equations based on the modified entropy. By doing so, we derive the modified Friedmann equations for the universe. This study, in addition to revealing the possible features and constraints of the model, holds the potential to introduce new methods for explaining observational data in cosmology.
@@ IV Deriving the Modified Friedmann Equations
To deepen our understanding of entropy in this context, we analyze the luminosity distance and its dependence on redshift, both of which are pivotal for studying cosmic dynamics. The luminosity distance, which quantifies how an object’s apparent brightness relates to its intrinsic luminosity, plays a central role in probing the universe’s expansion. It is a function of redshift ($z$), which measures the stretching of light caused by cosmic expansion. The relationship is expressed as:
@@
In cosmology, the distance modulus ($\mu$) is a widely used parameter, defined as $\mu\equiv m-M$, where $m$ represents the apparent brightness and $M$ is the absolute brightness of an object. Its relationship with the luminosity distance is expressed as
@@
To evaluate the deviation of the model from observational data and determine the best fit, we employed the chi-squared ($\chi^{2}$) parameter, defined as
@@ V Summary and Conclusion
Our analysis reveals that small corrections produce negligible deviations from standard cosmology, while moderate values yield significant changes in cosmological behavior. Larger corrections ($\gamma>0.47$) result in inconsistencies with observational constraints, including the distance modulus versus redshift data from the supernovae data set. This analysis of compatibility with empirical data demonstrates that quantum fluctuations can lead to a maximum $47\%$ increase in event horizon area without conflicting with observations. These findings highlight the sensitivity of the universe’s evolution to quantum-gravitational effects and offer an alternative framework to address cosmological problems. The study provides a quantitative basis for further exploration of quantum gravity influence on cosmology, with observational constraints on $\gamma$ serving as a critical bridge between theory and experiment. Future work could refine these constraints with different observational datasets and extend the framework to include additional quantum-gravitational phenomena. This investigation underlines the potential of quantum entropy corrections to address unresolved questions in fundamental physics and cosmology.

### v2

@@ II Cosmology of Holographic Entropies
However, the crucial issue is related to resolving the crisis faced by the standard model of cosmology. In this direction, akin to modified gravity theories, modifying the entropy associated with the horizon also influences the equations governing the evolution of the universe Nojiri et al. (2021). In fact, there are many proposals for generalized versions of the entropy. Taking into account systems with long-range interactions, such as gravitational systems, which are categorized as non-extensive systems, Tsallis introduced a generalization of standard thermodynamics in 1988 Tsallis (1988). This extension broadens its applicability to non-extensive scenarios while retaining the standard Boltzmann-Gibbs theory as a limiting case Tsallis (1988). It is, thus, noteworthy that, when equipped with a specific entropy expression within any gravity theory, it becomes possible to rephrase the Friedmann equations of the Friedmann-Robertson-Walker (FRW) universe, which are extracted from the first law of thermodynamics on the apparent horizon, and vice versa. Considering this procedure, recent developments have led to significant interest in utilizing non-extensive Tsallis entropy for the entropy of gravitational systems Saridakis et al. (2018); Tavayef et al. (2018); da Silva and Silva (2021). This interest extends to the thermodynamic interpretation of gravitational field equations and the consequential effects of the non-extensive parameter within cosmology. Notably, this can result in variations in the gravitational constant’s strength, which consequently impacts the energy density of the dark components of the universe. This modification can explain the late-time acceleration of the universe without invoking a dark energy component Tavayef et al. (2018); Zamora and Tsallis (2022).
@@ III Generalized Entropy in Our Model
In this section, we propose an alternative form of entropy inspired by considerations from quantum gravity.
According to general relativity, the horizon of a black hole is precisely defined by the Schwarzschild radius. However, from a quantum mechanical perspective, understanding the geometry of spacetime requires measurements involving tools such as rods and clocks. These measurements are subject to inherent uncertainties Garay (1995); Kempf et al. (1995). For instance, attempting to measure a distance with high precision, especially at very short scales, encounters limitations imposed by the Heisenberg uncertainty principle. Precisely determining the position of the measuring rod introduces significant uncertainty in its momentum, which in turn results in uncertainty in the rod’s energy. Since energy contributes to spacetime curvature through general relativity, this energy uncertainty induces distortions in spacetime, leading to a fluctuating geometry. Thus, measurements at extremely small scales inherently face precision limits due to the interplay of quantum mechanics and general relativity.
This phenomenon persists even in the weak-energy regime of quantum gravity. Any consistent theory that fully reconciles gravity with quantum mechanics must accommodate this fundamental argument. The interplay of quantum uncertainty, energy fluctuations, and spacetime curvature highlights a fundamental limitation in measurement precision, underscoring the need for a deeper understanding of quantum spacetime.
Figure 1: An exaggerated representation of a black hole’s surface at Planck-scale resolution, illustrating quantum fluctuations and the fuzziness of the event horizon.
Given the limitations on measurement precision imposed by quantum mechanics, it is reasonable to assume a degree of roughness in the surface area of a black hole when examined at the small scales, as shown in Fig. 1. This roughness arises as a random distribution, potentially rooted in a quantum theory of gravity where uncertainty plays a pivotal role. Larger surface areas correspond to more fluctuations, while smaller surfaces exhibit fewer fluctuations.
Assuming that these fluctuations are homogeneously distributed, the number of fluctuations scales proportionally with the surface area. For example, doubling the surface area results in a doubling of the number of fluctuations. Based on this proportionality, the corrected surface area can be expressed as:
@@
where $A_{0}$ represents the area of the black hole as defined by its Schwarzschild radius, and the second term accounts for the surface roughness. Here, $\gamma$ is a parameter quantifying the quantum mechanical fuzziness in the black hole’s surface area. While the average volume remains unchanged, tiny ripples can significantly increase the surface area and intuitively we expect $\gamma$ is always positive.
Building on this concept, the generalized form of black hole entropy can be written as:
@@
where $A_{0}$ again denotes the horizon area. To test the validity of this model, we examine its cosmological implications by employing the thermodynamic interpretation of the gravitational field equations, modified by the generalized entropy. This leads to the derivation of modified Friedmann equations for the universe. By exploring these modifications, we can uncover new features and constraints of the model and potentially introduce innovative methods for explaining observational data in cosmology.
@@ IV Deriving the Modified Friedmann Equations
To deepen our understanding of entropy in this context, we analyze the luminosity distance and its dependence on redshift, both of which are pivotal for studying cosmic dynamics. The luminosity distance, which quantifies how an object’s apparent brightness relates to its intrinsic luminosity, plays a central role in probing the universe’s expansion. It is a function of redshift ($z$), which measures the stretching of light caused by cosmic expansion. The relationship is expressed as Weinberg (2008)
@@
In cosmology, the distance modulus ($\mu$) is a widely used parameter, defined as $\mu\equiv m-M$, where $m$ represents the apparent brightness and $M$ is the absolute brightness of an object. Its relationship with the luminosity distance is expressed as Weinberg (2008)
@@
To evaluate the deviation of the model from observational data and determine the best fit, we employed the chi-squared ($\chi^{2}$) parameter, defined as Witkov and Zengel (2019)
@@ V Summary and Conclusion
Our analysis reveals that small corrections produce negligible deviations from standard cosmology, while moderate values yield significant changes in cosmological behavior. Larger corrections ($\gamma>0.47$) result in inconsistencies with observational constraints, including the distance modulus versus redshift data from the supernovae data set. This analysis of compatibility with empirical data demonstrates that quantum fluctuations can lead to a maximum $47\%$ increase in event horizon area without conflicting with observations. These findings highlight the sensitivity of the universe’s evolution to quantum-gravitational effects and offer an alternative framework to address cosmological problems. The study provides a quantitative basis for further exploration of quantum gravity influence on cosmology, with observational constraints on $\gamma$ serving as a critical bridge between theory and experiment.
Moreover, this model logically follows from the effects of quantum mechanics’ random distributions on spacetime geometry in the low-energy limit. Consequently, any quantum gravity theory consistent with quantum mechanical foundations is expected to exhibit similar properties.
Future work could refine these constraints with different observational datasets and extend the framework to include additional quantum-gravitational phenomena. This investigation underlines the potential of quantum entropy corrections to address unresolved questions in fundamental physics and cosmology.